### v1

# BioMoDiffuse: Physics-Guided Biomechanical Diffusion for Controllable and Authentic Human Motion Synthesis

Abstract
Human motion generation holds significant promise in fields such as animation, film production, and robotics. However, existing methods often fail to produce physically plausible movements that adhere to biomechanical principles. While recent autoregressive and diffusion models have improved visual quality, they frequently overlook essential biodynamic features, such as muscle activation patterns and joint coordination, leading to motions that either violate physical laws or lack controllability. This paper introduces BioMoDiffuse, a novel biomechanics-aware diffusion framework that addresses these limitations. It features three key innovations: (1) A lightweight biodynamic network that integrates muscle electromyography (EMG) signals and kinematic features with acceleration constraints, (2) A physics-guided diffusion process that incorporates real-time biomechanical verification via modified Euler-Lagrange equations, and (3) A decoupled control mechanism that allows independent regulation of motion speed and semantic context. We also propose a set of comprehensive evaluation protocols that combines traditional metrics (FID, R-precision, etc.) with new biomechanical criteria (smoothness, foot sliding, floating, etc.). Our approach bridges the gap between data-driven motion synthesis and biomechanical authenticity, establishing new benchmarks for physically accurate motion generation.
Figure 1: Our model, BioMoDiffuse, generates physically realistic motions based on textual input. By embedding Euler-Lagrange dynamics into the generation process, the model produces physical-optimized trajectories that inherently satisfy physical laws, achieving superior motion authenticity compared to current SOTA approaches. Specifically, BiMoDiffuse mitigates common kinematic artifacts including foot sliding, body interpenetration, and motion discontinuities while ensuring temporal smoothness.
1 Introduction
Human motion generation has wide-ranging applications in animation, film, and robotics. In recent years, there has been growing focus on generating human motion conditioned on language, sound, and motion trajectories [68, 61, 12, 49]. Among these, language-guided motion generation stands out due to its high practical value, complexity, and the extensive foundational research supporting the field. Recent advancements [15, 48, 66, 17, 22, 24, 50, 88] in autoregressive and diffusion models have significantly improved the synthesis of diverse motion sequences while maintaining semantic consistency.
While existing methods have made strides in enhancing the visual plausibility of motion sequences, studies in biomechanics and animation highlight significant shortcomings[54, 10] of them in capturing key aspects of human motion. The absence of explicit modeling for biomechanical and dynamic features prevents current models from learning essential information, such as acceleration patterns and muscle activation levels. Consequently, generated motions are often rigid and physically unrealistic, deviating significantly from natural human movement in rhythm and joint coordination. Additionally, current evaluation metrics fail to quantitatively assess physical plausibility and motion velocity, while existing methods struggle with precise control over movement amplitude and speed. These limitations considerably diminish the practical value of motion generation in professional applications that require biomechanical accuracy.
To address these challenges, we propose BioMoDiffuse, a biomechanical-aware motion generation framework that seamlessly integrates musculoskeletal dynamics with diffusion processes. Our key innovation lies in embedding explicit physical supervision into the diffusion pipeline, leveraging a physics-aware network based on modified Euler-Lagrange equations and physics guidance to enforce biomechanical plausibility. Specifically, BioMoDiffuse introduces real-time biomechanical verification during denoising and a decoupled motion control system for independent regulation of velocity and semantics. Additionally, since current evaluation protocols neglect biomechanical-related aspects, we establish a set of comprehensive evaluation metrics emphasizing physical realism, combining biomechanical validity metrics (e.g., motion smoothness, foot sliding rate) with traditional generation metrics (FID, Diversity). Experiments on HumanML3D [14] and KIT-ML [51] demonstrate significant improvements over existing methods.
The main contributions of this paper are as follows:
  • A biomechanics-guided diffusion framework that integrates EMG signals and kinematic features with acceleration-based physical supervision, ensuring compliance with rigid-body dynamics .
  • A physics-aware network architecture based on Euler-Lagrange principles, incorporating a novel motion control system for enhancing controllability.
  • A comprehensive evaluation framework that emphasizes physical realism, unifying biomechanical metrics and traditional metrics.
2 Related Work
Physical plausible motion processing. In motion generation, reinforcement learning (RL) has been utilized to fine-tune models for physical plausibility [72, 89, 67]. For example, ReinDiffuse [18] and PhysDiff [77] extend MDM [66] to improve motion quality. However, these RL-based approaches predominantly focus on refining foot-ground contact using opaque, black-box reward functions, often overlooking broader rigid-body dynamics and biomechanical principles. The lack of transparency in their loss design further constrains interpretability and limits precise control over kinematic behaviors. Alternatively, some methods incorporate auxiliary constraints during training [75, 29, 53, 34]; for instance, OmniControl [75] leverages spatial signals to guide generation. Nonetheless, these approaches tend to prioritize controllability at the expense of enforcing comprehensive physical laws, disproportionately emphasizing foot-centric adjustments rather than ensuring holistic motion plausibility.
Physical plausibility is also addressed in other motion-related tasks that demand biomechanically authentic synthesized motion sequences, such as human pose estimation from videos [32, 37, 33, 63, 60]. Recent advances try to incorporate physics-based modeling [78, 58, 87, 86, 59, 52, 35, 74] to enhance the quality of the estimated motion sequences. However, these methods still face challenges of suboptimal precision and excessive jitter. Although techniques like Gaussian filtering can suppress high-frequency noise, they often compromise precision. Consequently, several studies [76, 7, 25, 79, 69, 30, 65] have explored pose refinement techniques to reduce temporal instability. For instance, SmoothNet [79] employs a parametric approach to optimize the temporal dimension, substantially enhancing motion quality.
Diffusion models. Diffusion models have demonstrated remarkable success in image generation, with various methods leveraging language guidance to produce high-quality images [55, 57, 81]. Among these approaches, DDPM [20] has emerged as the most widely adopted sampler, achieving impressive results. Building on DDPM, DDIM [62] reduces the number of inference steps while enabling deterministic sampling by adjusting the variance. The introduction of Classifier Guidance (CG) [9] empowered diffusion models to handle conditional generation tasks effectively, and CFG [19] further streamlined this process by replacing CG, enhancing the efficiency of guidance integration during training. In this paper, we observe that incorporating additional domain knowledge parallels the mechanisms of CG and CFG conceptually, leading us to re-derive the diffusion process accordingly.
In recent years, many studies have tackled the Diffusion Inversion problem, which provides a method for mapping real images to latent space, with broad applications in image editing [46, 71, 36, 26, 85, 8]. However, Diffusion Inversion methods have rarely been applied to tasks outside of image generation. We observe that both motion refinement and image editing aim to preserve the macroscopic structure while modifying finer details. Consequently, we are the first to apply DDIM Inversion to the task of motion refinement.
Text-driven human motion generation. Early approaches to text-driven human motion generation conceptualized motion as a form of language, adopting autoregressive methods commonly used in language models [2, 1, 48, 3]. Recognizing the potential of large language models, subsequent studies improved motion embeddings to enable the joint learning of motion and text, yielding promising results [15, 16, 80, 11, 73, 38, 42].
MDM [66] pioneered the application of diffusion models to human motion generation, while MLD [5], inspired by VAE [31], introduced a diffusion process within the latent space. MotionDiffuse [82] established a larger model framework, using cross-attention to link motion and text. Building on MotionDiffuse, ReMoDiffuse [83] incorporated retrieval-based motion databases and supplementary knowledge, resulting in significant performance improvements. MoFusion [43] and StableMoFusion [22], drawing inspiration from Stable Diffusion [55], utilized the UNet [56] architecture without relying on external knowledge, achieving state-of-the-art results. Notably, experiments with StableMoFusion revealed that DiT [47] and RetNet [64] underperformed relative to UNet, validating our choice of UNet as the backbone for the diffusion framework.
3 Method
Current approaches fail to capture the essential characteristics of human motion due to the absence of biomechanical constraints and rigid-body dynamics modeling, limiting their ability to learn authentic movement patterns. To address this gap, we present a novel motion generation framework that explicitly incorporates biomechanical principles and rigid-body kinematics, which is capable of synthesizing authentic physically plausible human motions. Our methodological contributions are as follows: Sec. 3.1 establishes the theoretical foundation by formalizing the biomechanical constraints; Sec. 3.2 introduces a physics-aware network architecture derived from Euler-Lagrange equations, ensuring dynamic compliance; Sec. 3.3 presents our biomechanics-guided diffusion based on joint probability distribution analysis. The complete motion generation framework is detailed in Sec. 3.4. Furthermore, Sec. 3.5 demonstrates that our proposed method, equipped with diffusion inversion, can be effectively utilized for physics-aware motion refinement.
Figure 2: Overview of BioMoDiffuse. At each diffusion timestep $t$, our framework processes the motion sequence $\Theta^{t}$ alongside text conditions. The text is encoded using a frozen CLIP encoder and concatenated with $\Theta^{t}$ as input to the UNet backbone. The light brown panel (left) illustrates our physics-aware network, which consists of: (1) An Accel Net that implements augmented Euler-Lagrange equations to predict biomechanically valid accelerations, and (2) An Autoregressive Block that models temporal dependencies. The training objectives combine: $\mathcal{L}_{\text{phys}}$ (from Accel Net), $\mathcal{L}_{\text{diffusion}}$ (reconstruction), and $\mathcal{L}_{\text{accel}}$ (acceleration supervision against ground truth). Structural details of the orange (autoregressive block) and light blue (Accel Net) components are provided in the right panels.
3.1 Preliminaries
Text-guided human motion generation requires an input text to condition the generation of a motion sequence $\Theta\in\mathbb{R}^{T\times D}$, where $T$ denotes the sequence length and $D$ represents the dimension of the joint representation. This joint representation can be converted into a coordinate representation $\Phi\in\mathbb{R}^{T\times 3d}$, where $d$ is the number of joints. The Euler-Lagrange equation is given by:
$F=M\cdot\ddot{\Phi_{t}}+C,$ (1)
where $F$ represents the internal and external contact forces, $M$ is the inertia matrix, and $C$ captures the combined effect of the Coriolis force and gravity. The term $F$ can be decomposed into internal force $I$ and the external contact force $E$. According to biomechanical principles, the internal force $I$ is generated by muscle movements and can be expressed as:
$I=R\cdot A(u),$ (2)
where $R$ is the muscle rotation matrix, $A(u)$ is the muscle activation force, and $u$ is the muscle activation coefficient. Using this decomposition, the Euler-Lagrange equation for human motion can be rewritten as Eq. 3. Please refer to Appendix D for details.
$\ddot{\Phi_{t}}=M^{-1}(E+C+R\cdot A(u)).$ (3)
3.2 Physics-Aware Network
The Euler-Lagrange equation-based network provides a stronger prior than conventional neural networks, as it inherently captures acceleration dynamics while reducing problem complexity. A crucial design here is to directly embed the constraints within the network architecture itself. In a physical sense, the Euler-Lagrange equation serves as a rigorous constraint for human motion: by framing motion generation as $Motion=Gen(text)$, this equation ensures the learned function aligns closely with real-world biomechanics; From a deep learning perspective, Euler-Lagrange equation networks offer a more robust alternative to traditional non-parametric constraints for handling diverse human skeleton data in supervised settings. Additionally, the inherent properties of Euler-Lagrange equations in our problem context guarantee convergence to a stable and unique solution.
Following the MIA dataset representation, a learnable mapping exists between human poses and the activation coefficients of primary muscles $u\in\mathbb{R}^{8}$, defined as $u=u(\Phi_{:t})$. The muscle rotation matrix $R$ depends solely on pose rotation, while the inertia matrix $M$—a symmetric matrix—is intrinsic to the body and determined by position. The Coriolis force and gravity are formulated as:
$C=\dot{M}(\Phi_{t})\dot{\Phi_{t}}-\frac{1}{2}\dot{\Phi_{t}}^{T}\frac{\partial M}{\partial\Phi_{t}}\dot{\Phi_{t}}+G(\Phi_{t}),$ (4)
which are also pose-dependent. External contact forces $F$ (e.g., elastic/resistive forces) depend on both position and velocity. Thus, Eq. 3 is reformulated as:
$\ddot{\Phi}_{t}=M^{-1}(\Phi_{:t})\left(F(\Phi_{:t})+C(\Phi_{:t})+R(\Phi_{:t})\odot A(u)\right),$ (5)
where $\odot$ denotes element-wise multiplication. We implement three neural networks to approximate $M$, $F$, and $C$, with $M$ constrained as symmetric. The term $R\odot A(u)$ is modeled via cross-attention, enabling lightweight networks to predict pose acceleration. Finally, leveraging autoregressive processing, the predicted $\ddot{\Phi}_{t}$ and speed control signal are integrated as inputs to derive the optimized $\Theta_{t}$, ensuring temporal coherence and biomechanical fidelity.
3.3 Physics Guidance
We introduce additional guidance into the diffusion process to ensure that acceleration remains physically meaningful and achieve more stable training and faster convergence. Our network explicitly treats the acceleration $\ddot{\Phi_{t}}$ as an intermediate variable. We aim for the acceleration $\ddot{\Phi_{t}}$ in our physics-based network to converge to the real result as quickly as possible. To achieve this, we impose a supervised constraint on the acceleration during the diffusion process. Specifically, at each reverse diffusion step, we consider extending the distribution at step $t$, denoted by $\Theta^{t}$, into a joint distribution $S^{t}=\{\Theta^{t},\ddot{\Phi}^{t}\}$. According to DDPM theory, when performing diffusion on $S^{t}$, the model’s loss function can be written as:
$\min E_{Q(S^{0}:S^{T})}\left[-\log P_{\theta}(S^{0}|S^{1})\right.$ (6)
$+\sum_{t=2}^{T}\text{KL}(Q(S^{t-1}|S^{t},S^{0})\|\left.P_{\theta}(S^{t-1}|S^{t}))\right].$
We argue that, during the denoising process, the quantities $\Theta^{t-1},\ddot{\Phi}^{t-1}$ at step $t-1$ are unknown and do not satisfy the interpolation relationship. However, the known quantities $\Theta^{t},\ddot{\Phi}^{t}$ at step $t$ do satisfy this relationship. This leads to the following:
$\log P_{\theta}(S^{t-1}|S^{t})=\log P_{\theta}(\Theta^{t-1},\ddot{\Phi}^{t-1}|\Theta^{t},\ddot{\Phi}^{t})$ (7)
$=\log P_{\theta}(\Theta^{t-1},\ddot{\Phi}^{t-1}|\Theta^{t})$
$=\log P_{\theta}(\Theta^{t-1}|\ddot{\Phi}^{t-1},\Theta^{t})+\log P_{\theta}(\ddot{\Phi}^{t-1}|\Theta^{t}).$
In the expression in Eq. 7, $\log P_{\theta}(\Theta^{t-1}|\ddot{\Phi}^{t-1},\Theta^{t})$ resembles the condition $c$ in CG and CFG, while $\log P_{\theta}(\ddot{\Phi}^{t-1}|\Theta^{t})$ can be obtained using the physical network we propose. Thus, during training, we omit the physical network with probability $P_{c}$, and use it with probability $1-P_{c}$. We set $\ddot{\Phi}^{t-1}=\text{PhysNet}(\Theta^{t})$, and the predicted acceleration is compared to the ground truth acceleration, introducing a physics loss $L_{\text{phys}}=\text{MSE}(\text{PhysNet}(\Theta^{t}),\ddot{\Phi_{\text{gt}}}^{t})$. To further enhance acceleration accuracy, we design the following additional loss term for supervision:
$L_{\text{accel}}=\text{MSE}(\ddot{\Phi}^{t},\ddot{\Phi_{\text{gt}}}^{t}).$ (8)
Thus, the total loss function during training can be written as:
$L=L_{\text{diffusion}}+L_{\text{accel}}+L_{\text{phys}}.$ (9)
3.4 Model Architecture
Table 1: Quantitative evaluation on the HumanML3D and KIT-ML test set. $\pm$±indicates a 95% confidence interval. Bold face indicates the best result, while underscore refers to the second best. We reimplemented StableMoFusion based on the official code.
[…]
Top 1 Top 2 Top 3
KIT- ML MDM $0.164^{\pm.004}$ $0.291^{\pm.004}$ $0.396^{\pm.004}$ $0.497^{\pm.021}$ $9.191^{\pm.022}$ $10.85^{\pm.109}$ $1.91^{\pm.214}$
MLD $0.390^{\pm.008}$ $0.609^{\pm.008}$ $0.734^{\pm.027}$ $0.404^{\pm.027}$ $3.204^{\pm.027}$ $10.80^{\pm.117}$ $2.19^{\pm.071}$
T2M $0.370^{\pm.005}$ $0.569^{\pm.007}$ $0.693^{\pm.007}$ $2.770^{\pm.109}$ $3.401^{\pm.008}$ $10.91^{\pm.119}$ $1.48^{\pm.065}$
MotionDiffuse $0.417^{\pm.004}$ $0.621^{\pm.004}$ $0.739^{\pm.004}$ $1.954^{\pm.062}$ $2.958^{\pm.005}$ $11.10^{\pm.143}$ $0.73^{\pm.013}$
T2M-GPT $0.416^{\pm.006}$ $0.627^{\pm.006}$ $0.745^{\pm.006}$ $0.514^{\pm.029}$ $3.007^{\pm.023}$ $10.92^{\pm.108}$ $1.57^{\pm.039}$
ReMoDiffuse $0.427^{\pm.014}$ $0.641^{\pm.004}$ $0.765^{\pm.055}$ $0.155^{\pm.006}$ $2.814^{\pm.012}$ $10.80^{\pm.105}$ $1.24^{\pm.028}$
StableMoFusion $0.445^{\pm.006}$ $0.660^{\pm.005}$ $0.782^{\pm.004}$ $0.258^{\pm.029}$ $2.832^{\pm.014}$ $10.94^{\pm.077}$ $1.36^{\pm.062}$
Ours $0.448^{\pm.008}$ $0.666^{\pm.005}$ $0.788^{\pm.005}$ $0.211^{\pm.101}$ $2.772^{\pm.017}$ $11.11^{\pm.094}$ $1.38^{\pm.050}$
Human ML3D MDM $0.320^{\pm.005}$ $0.498^{\pm.004}$ $0.611^{\pm.007}$ $0.544^{\pm.044}$ $5.566^{\pm.027}$ $9.559^{\pm.086}$ $2.799^{\pm.072}$
MLD $0.481^{\pm.003}$ $0.673^{\pm.003}$ $0.772^{\pm.002}$ $0.473^{\pm.013}$ $3.196^{\pm.010}$ $9.724^{\pm.082}$ $2.413^{\pm.079}$
T2M $0.457^{\pm.002}$ $0.639^{\pm.003}$ $0.743^{\pm.003}$ $1.067^{\pm.002}$ $3.340^{\pm.008}$ $9.188^{\pm.002}$ $2.090^{\pm.083}$
MotionDiffuse $0.491^{\pm.001}$ $0.681^{\pm.001}$ $0.782^{\pm.001}$ $0.630^{\pm.001}$ $3.113^{\pm.001}$ $9.410^{\pm.049}$ $1.553^{\pm.042}$
T2M-GPT $0.491^{\pm.003}$ $0.680^{\pm.003}$ $0.775^{\pm.002}$ $0.116^{\pm.004}$ $3.118^{\pm.011}$ $9.761^{\pm.081}$ $1.856^{\pm.011}$
ReMoDiffuse $0.510^{\pm.005}$ $0.698^{\pm.006}$ $0.795^{\pm.004}$ $0.103^{\pm.004}$ $2.974^{\pm.016}$ $9.018^{\pm.075}$ $1.795^{\pm.043}$
StableMoFusion $0.545^{\pm.003}$ $0.743^{\pm.003}$ $0.833^{\pm.002}$ $0.133^{\pm.005}$ $2.814^{\pm.008}$ $9.704^{\pm.065}$ $1.844^{\pm.060}$
Ours $0.547^{\pm.003}$ $0.743^{\pm.002}$ $0.835^{\pm.002}$ $0.071^{\pm.003}$ $2.784^{\pm.008}$ $9.567^{\pm.086}$ $1.919^{\pm.063}$
Our biomechanics-aware diffusion framework is built upon three core components, as shown in Fig. 2. The architecture seamlessly integrates biomechanical priors with diffusion dynamics through:
Biomechanical Encoding. Using the MIA dataset [6], we pre-train a neuromuscular encoder that establishes a dynamical mapping $\mathcal{F}:\Theta\rightarrow u$ between joint angles $\Theta$ and muscle activation coefficients $u$. This encoder remains differentiable throughout end-to-end training.
Physics-Conditioned Diffusion. The UNet-based diffusion backbone processes motion sequences $\mathbf{x}_{t}$ at timestep $t$ through:
$z_{\theta}=\text{BioMoDiffuse}\big{(}\text{CLIP}(\mathbf{c}),\mathbf{x}_{t},\mathcal{F}(\Theta)\big{)},$ (10)
where text conditioning $\mathbf{c}$ is incorporated via cross-attention layers. We adopt a two-phase training process: initial diffusion followed by physics-constrained fine-tuning (which introduces the physics loss $\mathcal{L}_{\text{phys}}$).
Controllable Generation. Our decoupled control mechanism enables independent manipulation of motion dynamics and semantics through: (1) classifier-free guidance and physics guidance weights, and (2) the injection of biomechanical constraints via the pre-trained encoder.
3.5 Pose Refinement with Diffusion Inversion
Figure 3: Refining pose using our method. Building on existing pose estimation models, we integrate BioMoDiffuse with DDIM inversion [62]. By adopting an inversion-then-forward pipeline, our approach significantly improves the smoothness of generated SMPL sequences while enhancing the fitting accuracy of pose estimation.
Our proposed framework can be readily adapted to refine pose estimates from existing pose estimation methods. The key insight is that our physics-guided diffusion model is capable of generating physically plausible motions starting from latent noise. Given the initial estimates of human poses, we first employ diffusion inversion—a method that reverses the usual diffusion process—to recover the latent noise corresponding to the poses. We then feed this noise into the physics-guided diffusion model, which injects biomechanical constraints to produce the refined poses that are more authentic compared to original estimates. Specifically, given the initial SMPL [39] parameters $\mathbf{x}_{0}$ from pose estimation, we first invert them to timestep $t$ using DDIM sampling:
$\mathbf{x}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{0}+\sqrt{1-\alpha_{t}}\epsilon_{\theta}(\mathbf{x}_{0},t),$ (11)
where $t$ controls the balance between information preservation and noise introduction. The inverted latent $\mathbf{x}_{t}$ then undergoes forward diffusion with our physics guidance to produce the refined poses:
$\hat{\mathbf{x}_{0}}=f_{\theta}(\mathbf{x}_{t},t).$ (12)
Crucially, our parametric implementation inherits the physics constraints from Secs. 3.2 and 3.3 without the need for additional tuning, demonstrating a unified biomechanical framework across both motion generation and refinement tasks.
4 Experiments
4.1 Datasets and Metrics
We conducted training and testing on our model using both the KIT-ML [51] and HumanML3D [14] datasets. The KIT-ML dataset contains 3,911 motions and 6,363 natural language annotations, while HumanML3D is derived from and reannotates the HumanAct12 [13] and AMASS [44] datasets, comprising a total of 14,616 motion sequences and 44,970 descriptions. We follow the evaluation metrics used in previous motion generation works [66, 82, 22], including Frechet Inception Distance (FID), Precision, Diversity, Multimodality, and Multi-Modal Distance. Additionally, we consider physics-related metrics such as skating ratio, ground floating, ground penetration, and foot clipping, as used in prior studies [18, 75, 77]. Furthermore, we introduce two novel metrics based on Fourier analysis: Velocity Consistency (VC) and Acceleration Consistency (AC). Our motivation stems from the observation that while the Smoothness metric captures global temporal coherence, it overlooks the inherent kinematic relationship between stable body parts and agile extremities in human motion. For a motion sequence represented as $\Phi\in\mathbb{R}^{T\times 3d}$, the metrics are formally defined as:
VC $=\frac{1-\|\text{FFT}(\dot{\Phi})_{[:thre_{\text{low}}]}\|-\|\text{FFT}(\dot{\Phi})_{[thre_{\text{high}}:]}\|}{\|\text{FFT}(\dot{\Phi})_{[:thre_{\text{low}}]}\|+\|\text{FFT}(\dot{\Phi})_{[thre_{\text{high}}:]}\|},$ (13)
AC $=\frac{1-\|\text{FFT}(\ddot{\Phi})_{[:thre_{\text{low}}]}\|-\|\text{FFT}(\ddot{\Phi})_{[thre_{\text{high}}:]}\|}{\|\text{FFT}(\ddot{\Phi})_{[:thre_{\text{low}}]}\|+\|\text{FFT}(\ddot{\Phi})_{[thre_{\text{high}}:]}\|}.$
Spectral ratio analysis enables VC and AC to effectively characterize overall motion patterns by precisely quantifying the balance between high-frequency components (associated with rapid limb movements) and low-frequency components (correlated with stable torso dynamics). This frequency-domain approach provides a more accurate assessment of motion physicality compared to traditional smoothness metrics. Since the dataset includes motions like “climbing stairs” (a foot is floating from the ground) and “jumping while spinning” (abrupt acceleration), all these physics-based metrics should be made as consistent as possible with ground-truth instead of being blindly minimized to meet the characteristics of dataset.
4.2 Quantitative Results
Figure 4: User Study Results.
User Study.
We asked participants to rate four aspects of the generated motions on a 1–5 scale: “Generation Quality,” “Motion Smoothness,” “Physical Plausibility,” and “Text Alignment.” Using the HumanML3D test set. Thirty users were invited to compare the motion generation results of our model, StableMoFusion, and Ground Truth across 20 prompts. The results demonstrate that our model exhibits superior performance in terms of generation quality, motion smoothness, physical plausibility, and semantic alignment compared to StableMoFusion. Furthermore, on certain examples, our method achieves results that are comparable to the ground truth, showcasing its competitive performance.
Table 2: Quantitative evaluation of physical plausible metrics on the HumanML3D and KIT-ML test set. All experiments were repeated 20 times with different random seeds, and the mean results are reported in the tables. The results for T2M, and MotionDiffuse on HumanML3D is sourced from ReinDiffuse [18].
[…]
KIT- ML Real Motions 0.352 1.379 0.926 3.06 0.550 0.309 0.208 0.001
T2M $0.270$ $1.519$ $0.748$ $4.90$ $0.442$ $0.324$ $0.180$ $\bm{0.001}$
MotionDiffuse $0.258$ $1.861$ $1.177$ $1.92$ $0.136$ $0.808$ $0.490$ $0.005$
T2M-GPT $0.238$ $\bm{1.420}$ $0.832$ $3.97$ $0.503$ $0.267$ $0.188$ $0.002$
ReMoDiffuse $2.826$ $7.993$ $1.415$ $0.29$ $0.254$ $0.789$ $\bm{0.199}$ $0.006$
StableMoFusion $0.684$ $2.496$ $0.986$ $1.80$ $0.496$ $0.315$ $0.255$ $0.002$
Ours $\bm{0.426}$ $1.736$ $\bm{0.965}$ $\bm{2.25}$ $\bm{0.522}$ $\bm{0.306}$ $0.237$ $0.003$
Human ML3D Real Motions 0.124 0.682 0.358 2.60 0.205 0.057 0.000 0.000
T2M $0.326$ $2.826$ $0.301$ $1.13$ $0.110$ $0.217$ $0.235$ $0.263$
MotionDiffuse $0.082$ $0.963$ $0.332$ $1.17$ $0.031$ $0.426$ $0.386$ $0.217$
T2M-GPT $0.178$ $1.459$ $\bm{0.354}$ $1.28$ $0.192$ $0.052$ $0.022$ $\bm{0.001}$
ReMoDiffuse $2.353$ $7.163$ $0.605$ $0.15$ $0.040$ $0.376$ $0.019$ $0.008$
StableMoFusion $0.292$ $2.220$ $0.448$ $0.74$ $0.234$ $0.074$ $0.043$ $0.002$
Ours $\bm{0.135}$ $\bm{0.959}$ $0.385$ $\bm{1.45}$ $\bm{0.204}$ $\bm{0.056}$ $\bm{0.013}$ $0.002$
Main Results.
Tab. 1 and Tab. 2 compare our method with existing approaches across both physical plausibility metrics (e.g., smoothness, foot sliding) and traditional generation metrics (FID, Diversity). Our model achieves state-of-the-art results in biomechanical authenticity while simultaneously advancing traditional motion quality, outperforming prior works across multiple dimensions. Experiments reveal critical limitations in existing paradigms: GPT-style generative models tend to over-smooth motions, failing to capture realistic dynamics, while diffusion-based methods exhibit temporal discontinuity—likely due to their implicit treatment of spatiotemporal relationships without explicit temporal modeling. By integrating our physics-aware network into the diffusion framework, we address these issues: the network’s frame-wise acceleration recalculation enhances temporal dependency learning, directly improving physical metrics. Concurrently, the introduction of biomechanical priors reduces solution space complexity, enabling superior performance in traditional metrics without compromising motion diversity.
Figure 5: Qualitative comparison of the state-of-the-art methods. Quantitative evaluations demonstrate that BioMoDiffuse achieves superior language-motion alignment, enhanced temporal smoothness, and improved physical plausibility compared to current approaches.
4.3 Ablation Study
Table 3: Ablation Study on traditional metrics. we evaluate the impact of adding $L_{accel}$ loss, incorporating the Physical Network, and introducing Physics Guidance on the performance of our method. All experiments were repeated 20 times with different random seeds, and the mean results are reported in the tables.
[…]
Top 1 Top 2 Top 3
Ours $\times$ $\times$ $\times$ $0.090$ $2.824$ $9.590$ $1.849$ $0.542$ $0.739$ $0.832$
$\checkmark$ $\times$ $\times$ $0.129$ $2.835$ $\bm{9.659}$ $1.882$ $0.537$ $0.734$ $0.827$
$\checkmark$ $\checkmark$ $\times$ $\bm{0.066}$ $\bm{2.783}$ $9.596$ $\bm{1.948}$ $\bm{0.549}$ $\bm{0.746}$ $\bm{0.836}$
$\checkmark$ $\checkmark$ $\checkmark$ $0.071$ $2.784$ $9.567$ $1.919$ $0.547$ $0.743$ $0.835$
Table 4: Ablation Study on physical plausible metrics.
[…]
GT - - - $0.124$ $0.682$ $0.358$ $2.60$ $0.205$ $0.057$ $0.000$ $0.000$
Ours $\times$ $\times$ $\times$ $0.292$ $2.220$ $0.448$ $0.74$ $0.234$ $0.074$ $0.043$ $0.002$
$\checkmark$ $\times$ $\times$ $0.144$ $1.022$ $0.437$ $1.24$ $0.228$ $0.076$ $0.052$ $0.003$
$\checkmark$ $\checkmark$ $\times$ $0.135$ $\bm{0.954}$ $0.421$ $1.39$ $0.227$ $0.060$ $0.026$ $0.002$
$\checkmark$ $\checkmark$ $\checkmark$ $\bm{0.135}$ $0.959$ $\bm{0.385}$ $\bm{1.45}$ $\bm{0.204}$ $\bm{0.056}$ $\bm{0.013}$ $\bm{0.002}$
Our method introduces three key innovations to the diffusion framework: acceleration loss, a physics network grounded in Euler-Lagrange equations, and physics-guided sampling. Ablation studies under identical training settings Tab. 3 and Tab. 4 validate their contributions. While the acceleration loss enhances kinematic feature extraction by expanding the motion representation dimension from 263 to 329, it initially degrades generation quality (e.g., FID deteriorates by 43.3%) due to increased solution space complexity. The proposed physics network addresses this by embedding biomechanical constraints, reformulating motion generation as a constrained optimization problem. This reduces spatial complexity while enforcing rigid-body dynamics, leading to a 48.8% FID improvement with only a 6.68% parameter increase over the UNet baseline. Crucially, the network’s Euler-Lagrange foundation enables simultaneous gains in physical plausibility and generation quality. Finally, physics-guided sampling prioritizes biomechanical fidelity during denoising, stabilizing acceleration fitting and further improving physical metrics, albeit with a slight trade-off in generation scores—a phenomenon consistent with CFG principles.
4.4 Motion Refining Results
Table 5: Comparative results for pose refinement. We trained and tested the WHAM [60] baseline and WHAM $+$ our method using the official code on the 3DPW [70] and RICH [21] datasets. Bold numbers denote the most accurate method in each column. Accel is in $m/s^{2}$, all other errors are in $mm$.
[…]
3DPW WHAM $36.34$ $61.46$ $70.64$ $6.52$
+ Inversion $\bm{36.13}$ $\bm{61.11}$ $\bm{70.52}$ $\bm{5.55}$
RICH WHAM $50.67$ $83.00$ $95.32$ $5.35$
+ Inversion $\bm{50.60}$ $\bm{81.80}$ $\bm{94.84}$ $\bm{4.06}$
As shown in Tab. 5, we validate our model’s motion refinement capability using WHAM [60], a state-of-the-art pose estimation framework that leverages temporal modeling to achieve leading results on metrics like acceleration error. By integrating our physics-aware network with WHAM via diffusion inversion, we significantly improve motion smoothness while also advancing pose prediction accuracy to new state-of-the-art levels. This indicates that our model, equipped with diffusion inversion, can be applied after any pose estimation module to refine the motion sequence, resulting in more fluent and physically plausible results. These results demonstrate that our physical network effectively enhances temporal continuity and overall robustness, even without acceleration loss supervision. This dual capability underscores the versatility of our approach, showcasing its ability to refine both generated and estimated motions by enforcing physical plausibility at the intersection of kinematics and dynamics.
4.5 Ablation Experiments on Speed Control Signal
Table 6: Ablation Study on weight of Speed Control Signal. We evaluate the impact of different Speed Control Signal on generation quality and physical plausibility metrics. All experiments use the HumanML3D dataset.
[…]
GT $0.002$ $2.974$ $0.797$ $0.124$ $0.682$ $0.358$ $2.60$ $0.205$ $0.057$ $0.000$ $0.000$
$0.5$ $0.479$ $2.860$ $0.818$ $0.148$ $1.085$ $0.294$ $1.83$ $0.096$ $0.070$ $0.012$ $0.001$
$0.8$ $0.114$ $2.782$ $0.834$ $0.138$ $0.989$ $0.356$ $1.56$ $0.165$ $0.061$ $0.013$ $0.002$
$1.0$ $0.071$ $2.784$ $0.835$ $0.135$ $0.959$ $0.385$ $1.45$ $0.204$ $0.056$ $0.013$ $0.002$
$1.5$ $0.085$ $2.824$ $0.830$ $0.134$ $0.937$ $0.428$ $1.32$ $0.276$ $0.050$ $0.013$ $0.002$
$2.0$ $0.124$ $2.869$ $0.824$ $0.136$ $0.942$ $0.447$ $1.26$ $0.321$ $0.045$ $0.013$ $0.002$
Our physical network incorporates a speed control signal within the autoregressive block. Experiments reveal that applying this signal only during inference (via input weighting) dynamically scales motion amplitude, with stronger signals producing larger movements. However, three critical observations emerge: (1) Traditional generation metrics (e.g., FID) remain unaffected by signal strength, aligning with the convention that model parameters require consistency between training and testing; (2) Certain physical metrics, such as ground penetration, foot clipping, and velocity/acceleration consistency, also show minimal sensitivity; (3) Metrics tied to acceleration magnitude—smoothness and skating ratio—decrease with stronger signals, while ground floating increases. We attribute this to the explicit acceleration modeling in our framework: since the network learns absolute acceleration values (rather than relative relationships), scaling the input signal alters acceleration magnitudes, thereby directly impacting physics-based metrics. This demonstrates our model’s unique ability to modulate motion amplitude through acceleration control while preserving generation quality—a trade-off avoided by conventional methods. Furthermore, these findings guide targeted optimization of physical plausibility by tuning acceleration-related parameters.
5 Conclusion
In this paper, we present BioMoDiffuse, a motion generation model built upon a physical network derived from extended Euler-Lagrange equations. This model incorporates the physical network into pose estimation using DDIM Inversion. BioMoDiffuse combines physics guidance with a state-of-the-art UNet backbone architecture, while simultaneously learning from electromyographic (EMG) signals and acceleration data. Experimental results indicate that BioMoDiffuse surpasses previous models in both generation quality and physical plausibility. However, due to current dataset limitations, BioMoDiffuse has not yet been tested on higher-frame-rate datasets with more comprehensive annotations. We leave it for future exploration.
Supplementary Material
Video:
We have provided supplemental videos in our supplemental material. In the supplementary video, we demonstrate the following: 1) the differences between the human skeletons generated by our method and those by StableMoFusion [22], 2) the differences in the skinned human models generated by our method and StableMoFusion, and 3) the text prompts used to generate these motions. All generations were produced using the same code under identical random seed conditions (seed=0). We reconstructed the meshes following the rendering approach provided by MLD [5].
Code:
The code has also been included in the supplementary materials, and detailed execution instructions can be found in the accompanying README file.
Appendix A Evaluation Metrics
In the design of physical plausibility metrics, we refer to the design of the Skating Ratio in Gmd [29], as well as the designs for Floating, Ground Penetration, and Foot Clipping in ReinDiffuse [18], as well as the Speed and Smoothness metrics provided by Fischer [10]. The specific definitions of these metrics are as follows:
  • Skating Ratio: Foot skating refers to the situation where a person’s foot is in contact with the ground but still has velocity. Given a human motion $\Theta$, the Foot Skating Ratio for the $i$-th frame can be expressed as:
    $FS^{i}(\Theta)=\exp\left(-||\dot{\Theta^{i}_{ft}}\cdot f^{i}_{S}\cdot f^{i}_{V}\cdot f^{i}_{\bar{V}}||_{2}\right)$ (14)
    where $\dot{\Theta^{i}_{ft}}$, $f^{i}_{S}$, $f^{i}_{V}$, and $f^{i}_{\bar{V}}$ represent the foot velocity, the contact label, the instantaneous velocity threshold of the foot, and the average velocity threshold of the foot at the $i$-th frame, respectively.
  • Floating: Floating refers to a scenario where the lowest joint of the generated motion is not in contact with the ground. Specifically, the Floating for the $i$-th frame can be represented as:
    $Float^{i}(\Theta)=\exp\left(-||(\Theta^{i}_{h}-h_{ground})\cdot f^{i}_{F}||_{2}\right)$ (15)
    where $\Theta^{i}_{h}$ denotes the y-axis coordinate of the lowest point of the action at the $i$-th frame, $h_{ground}$ is a threshold hyperparameter, set to 5cm in practice, and $f^{i}_{F}$ is a flag that equals 1 when $\Theta^{i}_{h}>h_{ground}$ and 0 otherwise.
  • Ground Penetration: Ground Penetration refers to the situation where the lowest joint of the generated motion penetrates below the ground. The Ground Penetration for the $i$-th frame can be expressed as:
    $GP^{i}(\Theta)=\exp\left(-||(h_{ground}-\Theta^{i}_{h})\cdot f^{i}_{P}||_{2}\right)$ (16)
    where $f^{i}_{P}$ is a flag that equals 1 when $\Theta^{i}_{h}<h_{ground}$ and 0 otherwise.
  • Foot Clipping: Foot Clipping refers to a situation where the two feet are too close together, causing mesh penetration. The Foot Clipping for the $i$-th frame can be expressed as:
    $FC^{i}(\Theta)=\exp\left(-||(\Theta^{i}_{lf}-\Theta^{i}_{rf})\cdot f^{i}_{C}||_{2}\right)$ (17)
    where $\Theta^{i}_{lf}$ and $\Theta^{i}_{rf}$ are the joint coordinates of the left and right feet, and $f^{i}_{C}$ is a threshold function. $f^{i}_{C}=1$ when the distance between the two feet exceeds a threshold, which is set to 5cm in practice.
  • Speed: Speed can be used to calculate the amplitude of human motion. The calculation for Speed at the $i$-th frame is given by:
    $Speed^{i}(\Theta)=\frac{||\Theta^{i}||_{2}}{\Delta_{t}}$ (18)
    where $\Delta_{t}$ is the sampling frame rate of the dataset, which is 0.05 for HumanML3D and 0.08 for KIT-ML.
  • Smoothness: Smoothness can be used to measure the jittering of the model’s acceleration. Models with poor quality tend to exhibit poor Smoothness. The calculation for Smoothness at the $i$-th frame is given by:
    $Smooth^{i}(\Theta)=\frac{|\ddot{\Theta^{i}}-\ddot{\Theta^{i-1}}|}{\Delta_{t}}$ (19)
It is important to note that all of the aforementioned metrics should ideally be as close as possible to the dataset. Since the dataset includes motions such as crawling, climbing stairs, and freestyle swimming, metrics like Skating Ratio and Floating are not necessarily better when smaller. For the KIT-ML dataset [51], Ground Penetration and Clipping are also not zero, which may be related to the quality of the KIT-ML dataset.
Appendix B Implementation Details
For training, we employ DDPM [20] with 1,000 denoising steps ($T=1,000$) and linearly varying variances $\beta_{t}$ from 0.0001 to 0.02 in the forward process. The model is trained using AdamW [40] with an initial learning rate of 0.0002 and a weight decay of 0.01 for 50,000 iterations during pre-training, followed by 10,000 iterations of fine-tuning with a batch size of 64 on an NVIDIA A40 GPU. The learning rate is reduced by a factor of 0.9 every 5,000 steps. Gradient norms are clipped to a maximum value of 1 during gradient descent. For class-conditional generation (CFG) [19], text conditioning $c$ is set to $\emptyset$ for 10% of the samples, and the physical network is ignored with a probability of $P_{c}=10\%$. Additionally, we apply Exponential Moving Average (EMA) to smooth the parameter updates and reduce mutations and oscillations.
For inference, we utilize the SDE variant of the second-order DPM-Solver++ [41] with Karras sigmas [28] for 10-step sampling. The scale for text conditioning CFG is set to 2.5.
In our implementation of pose refinement, we used WHAM [60] as the baseline. However, the official WHAM code repository only provides versions trained on five datasets: AMASS [44], InstaVariety [27], MPI-INF-3DHP [45], Human3.6M [23], and 3DPW [70]. It does not include a version that incorporates the BEDLAM [4] dataset for training. As a result, the baseline results in our work differ from those presented in the WHAM paper.
Appendix C Network Architecture
Table 7: Model Architecture of BioMoDiffuse on dataset HumanML3D.
Model Module Details
BioMo Diffuse Physical Network (1-1)
• GRU layer
• AccelNet:
   - Motion MLP
   - Muscle MLP
   - Physics Decoder
   - Muscle Decoder
   - M_Net
   - F_Net
   - C_Net
   - Pose2EMG
Cond Unet1D (1-2)
• Time MLP
• Downsampling Blocks (×4) :
   - Conv1D
   - Cross Attention
   - Residual Blocks
• Mid Blocks (×2) :
• Upsampling Blocks (×4) :
   - Transposed Conv1D
   - Residual Blocks
Tab. 7 illustrates the network architecture of BioMoDiffuse trained on HumanML3D [14]. Compared to the baseline UNet, the addition of our physical network only increases the parameter count by $6.68\%$. However, this results in a significant improvement in the FID score, which decreases by $48.8\%$ (from 0.129 to 0.068). Other metrics also show notable improvements, demonstrating that the proposed physical network is both lightweight and effective.
Appendix D Derivation of Euler-Lagrange equation
The classical Euler-Lagrange equation is formulated as:
$\frac{d}{dt}\left(\frac{\partial L}{\partial\dot{q}}\right)-\frac{\partial L}{\partial q}=Q,$ (20)
where $L=T-V$ is the Lagrangian, $T$ denotes kinetic energy, $V$ represents potential energy, $q$ are generalized coordinates, and $Q$ encompasses non-conservative forces such as joint torques and external contact forces.
By modeling the human body as a rigid-body system, its kinetic energy can be expressed as:
$T=\frac{1}{2}\dot{q}^{T}M(q)\dot{q},$ (21)
where $M(q)$ is the inertia matrix. The potential energy $V(q)$ is defined by the gravitational field. Substituting $L=T-V$ into the Euler-Lagrange equation yields:
$\frac{d}{dt}\left(\frac{\partial T}{\partial\dot{q}}\right)-\frac{\partial T}{\partial q}+\frac{\partial V}{\partial q}=Q.$ (22)
Substituting the explicit form of $T$, we derive:
$M(q)\ddot{q}+\dot{M}(q)\dot{q}-\frac{1}{2}\dot{q}^{T}\frac{\partial M}{\partial q}\dot{q}+G(q)=Q,$ (23)
where $G(q)=\frac{\partial V}{\partial q}$. Leveraging the symmetry of $M(q)$, we reorganize indices and employ Christoffel symbols to define the Coriolis matrix:
$C(q,\dot{q})\dot{q}=\sum_{i,j}\frac{1}{2}\left(\frac{\partial M_{kj}}{\partial q_{i}}+\frac{\partial M_{ki}}{\partial q_{j}}-\frac{\partial M_{ij}}{\partial q_{k}}\right)\dot{q}_{i}\dot{q}_{j}.$ (24)
It can be rigorously shown that $C(q,\dot{q})\dot{q}=\dot{M}(q)\dot{q}-\frac{1}{2}\dot{q}^{T}\frac{\partial M}{\partial q}$, demonstrating that the Coriolis term arises from the derivative of the inertia matrix and quadratic velocity coupling, which depend solely on the body’s pose.
Figure 6: More samples from our BioMoDiffuse for text-to-motion synthesis. Samples generated with text prompts of HumanML3D’s test set. We recommend the supplemental video to see more motion results.
Appendix E Comparison on Samplers
Table 8: Comparative Evaluation of Motion Generation Methods. We compare different methods on generation quality, efficiency, and physical plausibility metrics using the HumanML3D dataset.
[…]
GT - - $0.002$ $2.974$ $0.797$ $0.124$ $0.682$ $0.358$ $2.60$ $0.205$ $0.057$ $0.000$ $0.000$
DEIS $10$ $3.659$ $0.148$ $2.732$ $0.843$ $0.298$ $2.388$ $0.380$ $0.91$ $0.195$ $0.054$ $0.013$ $0.002$
DEIS w/o Karras $10$ $3.615$ $0.146$ $2.723$ $0.844$ $0.271$ $2.285$ $0.381$ $0.88$ $0.195$ $0.051$ $0.013$ $0.002$
DPMSolver++ w/o Karras $10$ $3.664$ $0.126$ $2.755$ $0.840$ $0.147$ $1.071$ $0.374$ $1.39$ $0.196$ $0.050$ $0.012$ $0.001$
DPMSolver++ (20) $20$ $7.610$ $0.090$ $2.781$ $0.834$ $0.143$ $0.969$ $0.382$ $1.44$ $0.203$ $0.059$ $0.013$ $0.003$
DPMSolver++ (10) $10$ $2.553$ $0.071$ $2.784$ $0.835$ $0.135$ $0.959$ $0.385$ $1.45$ $0.204$ $0.056$ $0.013$ $0.002$
DDIM $50$ $19.82$ $0.123$ $2.741$ $0.839$ $0.132$ $0.933$ $0.368$ $1.57$ $0.185$ $0.054$ $0.012$ $0.002$
DDPM $1000$ $290.30$ $0.088$ $2.771$ $0.833$ $0.147$ $0.981$ $0.385$ $1.43$ $0.209$ $0.060$ $0.013$ $0.003$
In this section, we present the impact of different samplers on the generation performance of BioMoDiffuse. We evaluated the following samplers on our model: DDPM [20] , DDIM [62], DPMSolver [41], and DEIS [84]. Additionally, we incorporated Karras Sigma into DEIS and DPMSolver. Karras Sigma corresponds to constant-velocity thermal diffusion, which has been shown to ensure faster and higher-quality generation in the field of image synthesis. As shown in Tab. 8, we observe that DEIS fails to generate sufficiently smooth motions, while among the other samplers, DPMSolver achieves superior results in terms of FID and physical plausibility metrics. Furthermore, the inclusion of Karras Sigma contributes to improved FID scores. Based on these findings, we ultimately selected DPMSolver with Karras Sigma as the preferred configuration for our framework.
Appendix F User Study
Figure 7: User study interface for evaluating StableMoFusion and our model on HumanML3D.
In the user study, we generated multiple samples using the evaluation code on the HumanML3D test set, filtering out semantically redundant examples to create a final set of 20 distinct samples. Participants were asked to rate the generated motions on a 1–5 scale across four dimensions: ”Generation Quality,” ”Motion Smoothness,” ”Physical Plausibility,” and ”Text Alignment.” We invited 30 users to compare the motion generation results of our model, StableMoFusion, and Ground Truth based on these 20 prompts. To ensure unbiased feedback, the samples were presented in random order, preventing participants from identifying which motions were generated by our model.
Appendix G More Qualitative Results
Fig. 6 demonstrates the generation performance of our model across a broader set of prompts, while Fig. 8 provides a comparative analysis of our model against state-of-the-art methods on an extended set of prompts. The results indicate that our model consistently achieves superior performance across a wider range of motion types.
Figure 8: The comparison of the state-of-the-art methods on text-guided motion synthesis task. Samples generated with text prompts of HumanML3D’s test set. We recommend the supplemental video to see more motion results.